OpenStudy (rainbow_rocks03):

Solve for x. 2/3x = 10 A. 1/15 B. 6 2/3 C. 10 2/3 D. 15

TheSmartOne (thesmartone):

What is the reciprocal of 2/3 ?

OpenStudy (rainbow_rocks03):

IDK

TheSmartOne (thesmartone):

the reciprocal of \(\sf\Large\frac{a}{b}\) is \(\sf\Large\frac{b}{a}\)

TheSmartOne (thesmartone):

Now, can you tell me what the reciprocal of \(\Large\frac{2}{3}\) is @rainbow_rocks03

OpenStudy (rainbow_rocks03):

So the reciprocal is 3/2

OpenStudy (anonymous):

Dont just give the answer. Let @TheSmartOne help her. If you just give the answer, the person learned nothing

TheSmartOne (thesmartone):

correct, and when you multiply reciprocals you will get 1.

TheSmartOne (thesmartone):

So then, if we multipled the reciprocal on both sides of the equation. What would we get?